A white drawf star has a mass 10^30 kg. It's luminosity is 10^24j/s. Calculate how long it can survive with its present luminosity if it's internal temperature is 10^7k .
1
Expert's answer
2020-04-06T08:56:23-0400

The mass measured in solar masses:


"M=0.5M_\u2609."

The luminosity measured in solar luminosities:


"L=\frac{10^{24}}{3.828\cdot10^{26}}L_☉=0.0026L_☉."

The lifetime measured in solar lifetimes is:


"t=\frac{M}{L}=\frac{0.5M_☉}{0.0026L_☉}=192t_☉,"

or


"t=1.92\cdot10^{12}\text{ years}."